OpenStudy (anonymous):

Find an equation of a line through (2, 1) and perpendicular to −2x + 4y = 8.

OpenStudy (aripotta):

get −2x + 4y = 8 into slope-intercept form first

OpenStudy (anonymous):

so that's y=1/2x+2 right @AriPotta

OpenStudy (aripotta):

yes. so what is our slope?

OpenStudy (givemeyourpizza):

the slope has to be a negative reciprocal

OpenStudy (aripotta):

yea, since we want the equation that is perpendicular to that line, the slope needs to be the negative reciprocal of the given line's slope

OpenStudy (anonymous):

so -2

OpenStudy (aripotta):

yes :)

OpenStudy (givemeyourpizza):

if the slope is 1/2, then the negative reciprocal would be -2

OpenStudy (givemeyourpizza):

yes correct

OpenStudy (aripotta):

now that we have the slope and a point, we can plug that info into point-slope form, which is y - y1 = m(x - x1)

OpenStudy (aripotta):

y1 is the y part of the coordinate, and x1 is the x part of the coordinate. and m is the slope

OpenStudy (givemeyourpizza):

it would be y - 2 = -2(x - 1). I think.

OpenStudy (aripotta):

no

OpenStudy (aripotta):

you have it flip-flopped, @GiveMeYourPizza

OpenStudy (anonymous):

y-1=m(x-2) ?

OpenStudy (aripotta):

yes, but we know m. m is the slope, which we just found out was -2

OpenStudy (aripotta):

so y - 1 = -2(x - 2)

OpenStudy (givemeyourpizza):

oops...lol

OpenStudy (aripotta):

so that's your answer, unless it needs to be in slope-intercept form or standard form. if so, then you need to convert